OpenStudy (pugtato):

I need help understanding direct linear variation. Which does not show a direct variation between x and y? A. y=5x B. y=6/x C. y=0.7x D. y=x/9

OpenStudy (pugtato):

@Directrix @mathmale

jimthompson5910 (jim_thompson5910):

direct variation is in the form y = k*x in words (some variable) = (some constant) times (some other variable)

jimthompson5910 (jim_thompson5910):

so some examples would be y = 2*x y = 10*x y = 5.3*x y = (1/2)*x notice how they all fit the form y = k*x.

jimthompson5910 (jim_thompson5910):

does that make sense?

OpenStudy (pugtato):

so would the answer be D? since the constant and variable are flipped?

jimthompson5910 (jim_thompson5910):

D is a trick thrown at you y=x/9 is the same as y = (1/9)*x it's similar to saying (1/2)*3 = 3/2

jimthompson5910 (jim_thompson5910):

y = (1/9)*x is in the form y = k*x where k = 1/9

OpenStudy (pugtato):

Ohhh so in that case, would it be B?

jimthompson5910 (jim_thompson5910):

yes because y = 6/x cannot be written into the form y = k*x it's actually in the form y = k/x which is inverse variation

OpenStudy (pugtato):

Okay great, I think I got it now. Thank you so much! i'll tag you if I need more help with direct variation.

jimthompson5910 (jim_thompson5910):

you're welcome